A bottle heated with its mouth open from 15 degree Celsius to 100 degree Celsius .What fraction Of air originally contained in the vessel expelled

Grade:11

1 Answers

Vikas TU
7 years ago
From PV = nRT 
moles are inversely proportional to the temperature.
Therfore, n1T1= n2T2
or
n1*288 = n2*373
n1= 373*n2/288 => 1.2951*n2
or
n2 = 0.77214*n1
hence this would be the moles gets exposed when temperature reached to 373 K.